Grade 11General Physics

A bullet with a wooden plank of 1/20 velocity then then how. Many planks needed

7731971866
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
Expect that d be the thickness of the board and a be the speeding up of the board 
 
From the newton's third condition of movement 
 
v2 = u2+ 2ad 
 
On the off chance that we need to stop the shot then the last speed will be zero. 
 
Furthermore, on the off chance that we utilized n boards to stop that shot, 
 
That is, 
 
0 = u2 +2nad 
 
2nad = u2 
 
n = u2/(- 2ad) - (1.) 
 
As given in the inquiry, the speed is 1/20 speed 
 
that is v = u – u/20 
 
= 19u/20 
 
Substitute it in newtons third condition 
 
(19u/20)2 = u2 + 2ad 
 
\361/400 u^2 = u^2 + 2ad 
 
- 2ad = u^2(1 - 361/400) 
 
- 2ad = u^2 * 39/400 
 
Presently we need to substitute this incentive in condition (1.) 
 
n = u^2/(u^2 * 39/400) = 400/39 
 
Least number of boards = 400/39 = 11.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)